PLEASE help me with this!
A group of students is going on a field trip. If the group takes 3 vans and 1 car, 22 students can be transported. If the group takes 2 vans and 4 cars, 28 students can be transported. How many students can fit in each van?
I have come up with the equation 3v+1c=22 and 2v+4c=28. Im not sure if these are correct? PLEASE help! Thank you!

Answer by nerdybill(7384)

A group of students is going on a field trip. If the group takes 3 vans and 1 car, 22 students can be transported. If the group takes 2 vans and 4 cars, 28 students can be transported. How many students can fit in each van?
I have come up with the equation
3v+1c=22
2v+4c=28
What you have above is correct.
To solve we can apply the "elimination method"
Multiply both sides of the top equation by -4 and combine both equations:
-4(3v+1c)=-4(22)
2v+4c=28
.
-12v-4c=-88
2v+4c=28
------------
-10v = -60
v = 5 students can fit in each van